OpenStudy (anonymous):

Write 4 radical expressions that are each equivalent to 8 times the square root of x . One expression must use addition, one must use subtraction, one must use multiplication, and one must use division. Make sure that one of the expressions you create involves variables. PLEASEEEE HELPP!!

OpenStudy (anonymous):

Addition: 5√x + 3√x Subtraction: 10√x - 2√x Multiplication: 2√x * 4√x Division: 16√x / 2√x

OpenStudy (anonymous):

The multiplication and division is wrong

OpenStudy (anonymous):

multi- 2 ( x)).4 ( x) = 8 [( x)] = 8 x divis-4 ( x)) / 3 ( x)) = 8 x

OpenStudy (anonymous):

Teacher said thats still incorrect, can anyone please help?

OpenStudy (anonymous):

here is one \[\sqrt{64x}\]

OpenStudy (anonymous):

oh i see, multiplication would be \[4\sqrt{x}\times 2\] could work

OpenStudy (anonymous):

division you could use \[\frac{16\sqrt{x}}{2}\]

OpenStudy (anonymous):

Alright thank you i asked my teacher i'm waiting for a response